(xz)dx+(zy)dy=(x^2+y^2)dz
1
Expert's answer
2020-09-16T19:31:28-0400

Given differential equation is

"xzdx+zydy=(x^2+y^2)dz"


"z(xdx+ydy)=(x^2+y^2)dz"


"\frac{xdx+ydy}{x^2+y^2}=\frac{dz}{z}" .....(1)


let "x^2+y^2=t"

2xdx+2ydy=dt


Substitute in equation 1

"\frac{1}{2}\frac{dt}{t}=\frac{dz}{z}"

Integrating Both the side

"\frac{1}{2}\int\frac{dt}{t}=\int\frac{dz}{z}"


"\frac{1}{2}logt=logz+logc"


logt"^\frac{1}{2}=logzc"

"t^\frac{1}{2}=zc"



"(x^2+y^2)^\frac{1}{2}=zc"